OpenStudy (emogirl100):

A store had 235 MP3 players in the month of January. Every month, 30% of the MP3 players were sold and 50 new MP3 players were stocked in the store. Which recursive function best represents the number of MP3 players in the store f(n) after n months? f(n) = 0.7 × f(n − 1) + 50, f(0) = 235, n > 0 f(n) = 237 − 0.7 × f(n − 1) + 50, f(0) = 235, n > 0 f(n) = 0.3 × f(n − 1) + 50, f(0) = 235, n > 0 f(n) = 237 + 0.7 × f(n − 1) + 50, f(0) = 235, n > 0

OpenStudy (emogirl100):

@imqwerty @

OpenStudy (emogirl100):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

look for the pattern

OpenStudy (emogirl100):

?

OpenStudy (xapproachesinfinity):

month of February we have 253 lets call that first month okay

OpenStudy (emogirl100):

ok

OpenStudy (xapproachesinfinity):

so month 1 we have 253 month 2 we 0.3x253+50 month 3 we have 0.3x(0.3x253+50)+50

OpenStudy (xapproachesinfinity):

lets see the pattern after n months

OpenStudy (emogirl100):

kk

OpenStudy (xapproachesinfinity):

or you could just check if those options work

OpenStudy (xapproachesinfinity):

check for first month n=1

OpenStudy (xapproachesinfinity):

oh they made n=0 the first month not n=1 let's check for n=0 do we get 253 for what option

OpenStudy (emogirl100):

A?

OpenStudy (xapproachesinfinity):

oh well all of them actually have f(0)=253 for n=0 do n=1 and check

OpenStudy (emogirl100):

ok giv a sec

OpenStudy (xapproachesinfinity):

the right one has to have for n=1 f(1)=125.9

OpenStudy (xapproachesinfinity):

oh no error i used 253 instead of 235

OpenStudy (xapproachesinfinity):

for n=1 we should get f(1)=120.5 not 125.9 that was a mistake ok

OpenStudy (emogirl100):

ok

OpenStudy (xapproachesinfinity):

A and B or not good since they don't work out

OpenStudy (xapproachesinfinity):

same for D if you cheched

OpenStudy (xapproachesinfinity):

so C works as an answer

OpenStudy (emogirl100):

ok thank you

OpenStudy (xapproachesinfinity):

if you tried the find the pattern you would have gotten the same formula doing month 0 we have 235 month 1 we have 0.3x235+50 month 2 we have 0.3x(0.3x235+50)+50 and so on....

OpenStudy (xapproachesinfinity):

f(0)=235 f(1)=0.3x235+50=0.3xf(0) f(2)=0.3x(0.3x235+50)+50=0.3xf(1)+50 f(3)=0.3xf(2)+50 we go following the pattern tell n f(n)=0.3xf(n-1)+50

OpenStudy (xapproachesinfinity):

@emogirl100